How do I solve a problem that says "If Sonias Skis are $25 less than 3 times her bindings , and she spent $335 in total. How much were the skis
Answer by rfer(16322)

x=bindings
3x-25=skis
x+3x-25=335
4x=335+25
x=360/4
x=90 bindings
3x-25=245 skis